# CustomVideo: Customizing Text-to-Video Generation with Multiple Subjects

### v1

@@ Abstract
Customized text-to-video generation aims to generate high-quality videos guided by text prompts and subject references. Current approaches designed for single subjects suffer from tackling multiple subjects, which is a more challenging and practical scenario. In this work, we aim to promote multi-subject guided text-to-video customization. We propose CustomVideo, a novel framework that can generate identity-preserving videos with the guidance of multiple subjects. To be specific, firstly, we encourage the co-occurrence of multiple subjects via composing them in a single image. Further, upon a basic text-to-video diffusion model, we design a simple yet effective attention control strategy to disentangle different subjects in the latent space of diffusion model. Moreover, to help the model focus on the specific object area, we segment the object from given reference images and provide a corresponding object mask for attention learning. Also, we collect a multi-subject text-to-video generation dataset as a comprehensive benchmark, with 69 individual subjects and 57 meaningful pairs. Extensive qualitative, quantitative, and user study results demonstrate the superiority of our method, compared with the previous state-of-the-art approaches.
Figure 1: Customized text-to-video generation results of our proposed CustomVideo with given multiple subjects (left) and text prompts (below). Our approach can disentangle highly similar subjects, e.g., cat v.s. dog, preserving the fidelity of subjects and smooth motions.
@@ 1 Introduction
Text-to-video (T2V) generation Wang et al. (2023a); Zeng et al. (2023); Blattmann et al. (2023a); Kondratyuk et al. (2023) has achieved fantastic progress taking advantages of diffusion models Ho et al. (2020); Song et al. (2020); Lu et al. (2022). Recently, artists have dreamed of generating videos with their own belongings, e.g., pets, which directs a new research direction named customized T2V generation. Although existing methods Zhao et al. (2023); Wei et al. (2023); Jiang et al. (2023) have been proposed to generate videos from a single object, tackling multiple objects still remains a difficult scenario. The key challenge is to ensure the co-occurrence of multiple objects in the generated video and retain their corresponding identities.
@@
A recent work, VideoDreamer Chen et al. (2023b), proposes disen-mix finetuning and human-in-the-loop re-finetuning strategies based on Stable Diffusion Rombach et al. (2022), aiming to generate videos from multiple subjects. However, VideoDreamer falls short in guaranteeing the co-occurrence of multiple subjects and disentangling similar subjects due to its inconsistent object mixing strategy. As shown in Figure 3, the cat can not be consistently generated across all frames during inference. Moreover, VideoDreamer can not tackle a scenario with a foreground object with a background scene, such as car and barn (see Figure 3).
@@
In contrast, in our approach, we ensure the co-occurrence of multiple objects during model training, which encourages the model to capture the presence of different subjects simultaneously, thereby facilitating co-occurrence during inference. Additionally, we propose an attention control mechanism to disentangle multiple subjects during training, effectively guiding the model to focus on the corresponding subject area while disregarding irrelevant parts of the image. To facilitate this process, we incorporate a ground truth object mask, obtained through segmentation either from a model like SAM Kirillov et al. (2023) or provided by human annotators, as supervision during optimization. Our attention mechanism consists of two key designs that contribute to the disentanglement of subjects. Firstly, we highlight the corresponding subject area using the ground truth object mask on the cross-attention map, aligning learnable text tokens with subject identity. Secondly, we optimize the cross-attention map towards a slight negative value, excluding the desired subject and mitigating the influence of irrelevant areas in the input image. To comprehensively evaluate our proposed approach, we have curated a diverse dataset covering a wide range of categories including pet, person, toy, transportation, scene, plant, wearable item, and so on. Beyond the MultiStudioBench Chen et al. (2023b) which includes 12 subject pairs, our dataset CustomStudio comprises 69 individual subjects and 57 subject pairs, notably featuring some challenging scenarios involving visually similar objects. Through extensive experiments on this benchmark dataset, we provide qualitative, quantitative, and user study results that demonstrate the superiority of our method in generating high-quality videos with customized subjects. In summary, our contributions are as follows:
@@
  • We collect a multi-subject T2V dataset and build it as a comprehensive benchmark. We highlight that our benchmark covers a wide range of subject categories and diverse subject pairs over them.
  • Our method consistently outperforms previous state-of-the-art approaches in terms of diverse metrics, including Textual Alignment, Image Aligenment, and Temporal Consistency. More critically, in the user study, we obtain significant improvement with the scores 3×3$\times$3 × higher than these methods.
Figure 2: The overview of our proposed CustomVideo. We propose a simple yet effective co-occurrence and attention control mechanism with mask guidance to preserve the the fidelity of subjects for multi-subject driven text-to-video generation. During the training stage, only the key and value weights in the cross attention layers are fine-tuned. In the inference stage, given a text prompt integrated with learned text token, we can easily obtain high-quality videos with specific subjects.
@@ 2 Related Work
2.1 Text-to-Video Generation
@@
Text-to-video generation has made significant advancements in recent years Chen et al. (2023a); Duan et al. (2023); Wang et al. (2023a). Early approaches employed GANs Balaji et al. (2019); Skorokhodov et al. (2022) and VQVAE Hong et al. (2022); Villegas et al. (2023), while more recent works have explored diffusion models to generate high-quality videos He et al. (2022); Wang et al. (2023c); Zhang et al. (2023a, b). Make-A-Video Singer et al. (2023) utilizes a pre-trained image diffusion model with additional temporal attention finetuning. VideoLDM Blattmann et al. (2023b) introduces a multi-stage alignment approach in the latent space to generate high-resolution and temporally consistent videos. Other methods Khachatryan et al. (2023); Guo et al. (2023) generate videos with an image as the first frame and randomly initialized subsequent frames. To enhance controllability, VideoComposer Wang et al. (2023b) incorporates additional guidance signals, such as depth maps, to produce desired videos alongside text inputs. Tune-A-Video Wu et al. (2023) proposes a temporal self-attention module that fine-tunes a pre-trained image diffusion model, achieving successful generation of videos with specific text guidance. Furthermore, diffusion-based video-to-video editing approaches Qi et al. (2023); Geyer et al. (2023) have also been proposed for practical usages.
2.2 Subject-driven Customization
@@
There has been a growing interest in customizing pre-trained image and video diffusion models for personalized generation. Customization involves generating images and videos with specific subjects, typically based on a few reference images. For image diffusion customization, Textual Inversion Gal et al. (2023) represents a specific object as a learnable text token using only a few reference images. This learned text token can then be integrated into a sentence to generate personalized images during the inference stage. Additionally, DreamBooth Ruiz et al. (2023) fine-tunes the weights of the diffusion model to improve fidelity in image generation. Several works Kumari et al. (2023); Liu et al. (2023a, b) have explored personalized image diffusion with multiple subjects, focusing on parameter-efficient finetuning and text embedding learning. While there have been initial attempts to customize video diffusion, such as VideoAssembler Zhao et al. (2023) and VideoBooth Jiang et al. (2023), which use reference images to personalize the video diffusion model while preserving subject identity, and DreamVideo Wei et al. (2023), which decouples the learning process for subject and motion customization, these methods are designed for single objects and cannot handle multiple subjects when given. A recent work, VideoDreamer Chen et al. (2023b), proposes multi-subject driven video customization through disen-mix finetuning strategy with LoRA Hu et al. (2022). However, the generated videos do not guarantee the co-occurrence of multiple subjects, or disentanglement of different subjects. In this work, we propose a simple yet effective co-occurrence and attention mechanism that disentangles multiple subjects using masks as guidance while preserving the co-occurrence of subjects in the generated videos.
@@ 3.1 Preliminary: Text-to-Video
Video diffusion models (VDMs) Wang et al. (2023a); Duan et al. (2023); Chen et al. (2023a) generate videos by gradually denoising a randomly sampled Gaussian noise $\epsilon$, following an iterative denoising process that resembles a reverse procedure of a fixed-length Markov Chain. This iterative denoising allows VDMs to capture the temporal dependencies presented in video data. To be specific, a video diffusion model $\theta$ predicts the added noise at each timestep $t$ given a text condition $c$, where $t\in\{1,2,\cdots,T\}$. Thus, the training objective for this process can be expressed as a reconstruction loss:
@@
where $\boldsymbol{z}\in\mathbb{R}^{B\times L\times H\times W\times D}$ is the latent code of the input video with batch size $B$, video length $L$, height $H$, width $W$, and latent dimension $D$. $\epsilon_{\theta}$ is the noise prediction from the model. $\mathcal{T}$ is a pre-trained text encoder. $\boldsymbol{z}_{t}$ is obtained by adding noise to the ground truth $\boldsymbol{z}_{0}$ with $\boldsymbol{z}_{t}=\alpha_{t}\boldsymbol{z}_{0}+\sqrt{1-\alpha_{t}^{2}}\epsilon$, where $\alpha_{t}$ is a diffusion hyperparameter. In this work, we utilize zeroscope [^1] https://huggingface.co/cerspense/zeroscope_v2_576w T2V model as our base model, which is built upon a 3D UNet, with spatial and temporal modeling for generating high-quality videos.
@@ 3.2 CustomVideo with Multiple Subjects
Figure 2 provides an overview of our proposed CustomVideo framework, which enables multi-subject driven T2V customization. During the training stage, we employ a concatenation technique to ensure the co-occurrence of multiple subjects. Specifically, we combine these subjects into a single image, facilitating the learning of multi-subject patterns by the model. To address the challenge of entanglement among highly similar subjects, we propose an attention control mechanism. This mechanism ensures that the learnable word tokens align with their corresponding regions on the cross-attention map. By achieving this alignment, we enable the model to disentangle different subjects and enhance the quality of the generated videos. The training process focuses on training the subject-related learnable word tokens, as well as the key and value weights within the cross-attention layers of the UNet architecture, adopting a parameter-efficient finetuning approach. During the inference stage, users only need to provide a text prompt description integrated with the corresponding learned word tokens to generate high-quality videos aligned with their preferences. In the following, we will delve into the detailed workings of CustomVideo.
@@
In the context of multi-subject driven T2V generation, ensuring that the model consistently learns to generate videos with multiple subjects is crucial. In a previous work, VideoDreamer Chen et al. (2023b), the authors proposed a disen-mix strategy for multi-subject driven generation by fine-tuning the model using both single subject images and concatenated images of multiple subjects. However, we have identified that this mixing strategy can confuse the model due to the inconsistent number of subjects present in a single image. Consequently, it leads to unstable generation of videos with multiple subjects (see line ‘both single and concat’ in Figure 5). In our approach, we have found that fine-tuning the model using only concatenated images of multiple subjects is sufficient to ensure the co-occurrence of multiple subjects in the generated videos. Additionally, we have observed that providing clear subjects without background aids the model in learning the specific characteristics of the subjects more effectively. To achieve this, we perform background removal on the subject images. This can be accomplished manually or by employing an automatic tool such as the SAM model Kirillov et al. (2023).
@@
Ensuring the co-occurrence of multiple subjects in generated videos is achieved through image concatenation. However, a more challenging task is preserving the distinct characteristics of each subject without interference. Simply fine-tuning the model with concatenated images can lead to confusion among subject characteristics. As shown in Figure 5 (line ‘w/o pos. attn.’), the generated cat predominantly resembles the shape of the provided dog sample, despite having a similar color texture to the provided cat. Thus, disentangling the multiple subjects becomes crucial to generate high-quality videos that faithfully represent each subject’s characteristics.
@@ 3.3 Model Training and Inference
During training, we employ a parameter efficient fine-tuning strategy by only making the weights of key and value in all of the cross attention layers trainable. Inspired by previous text-to-image personalization Ruiz et al. (2023); Kumari et al. (2023), we conduct class-specific prior preservation to improve the diversity of generated videos and alleviate the issue of language drift. The loss for prior preservation is formulated as the following:
@@
where $\boldsymbol{z}^{\prime}$ is the latent code of the input class image. $\epsilon_{\theta}^{\prime}$ is the noise prediction from the model $\theta$ and $\epsilon^{\prime}$ is the randomly sampled Gaussian noise. $c^{\prime}$ is the text condition for class image and $t^{\prime}$ is the sampling timestep. To this end, we train our CustomVideo via an end-to-end manner, with the following overall training objective:
@@
During inference, CustomVideo only requires a specific text prompt with corresponding learned word token integrated to generate a required video. To be note that ground truth masks of subjects are not required in inference. With only text prompt as input, our proposed CustomVideo can generate videos with high diversity while not losing the the fidelity of subjects and motion smoothness.
@@ 4 Experiment
Figure 3: Qualitative results of our CustomVideo with comparison to SOTA methods, including DreamBooth, CustomDiffusion, and VideoDreamer. The first line indicates the given subjects, while each line indicates the frames generated by the corresponding method. The bottom line is the text prompt used during inference. We can observe that our CustomVideo can generate videos with much better fidelity of subjects compared with previous SOTA methods.

### v2

@@ Abstract
Customized text-to-video generation aims to generate high-quality videos guided by text prompts and subject references. Current approaches for personalizing text-to-video generation suffer from tackling multiple subjects, which is a more challenging and practical scenario. In this work, our aim is to promote multi-subject guided text-to-video customization. We propose CustomVideo, a novel framework that can generate identity-preserving videos with the guidance of multiple subjects. To be specific, firstly, we encourage the co-occurrence of multiple subjects via composing them in a single image. Further, upon a basic text-to-video diffusion model, we design a simple yet effective attention control strategy to disentangle different subjects in the latent space of diffusion model. Moreover, to help the model focus on the specific area of the object, we segment the object from given reference images and provide a corresponding object mask for attention learning. Also, we collect a multi-subject text-to-video generation dataset as a comprehensive benchmark, with 63 individual subjects from 13 different categories and 68 meaningful pairs. Extensive qualitative, quantitative, and user study results demonstrate the superiority of our method compared to previous state-of-the-art approaches. The project page is https://kyfafyd.wang/projects/customvideo.
Figure 1: Customized text-to-video generation results of our proposed CustomVideo given multiple subjects (left) and text prompts (below). Our approach can disentangle highly similar subjects, e.g., cat v.s. dog, simultaneously preserving the fidelity of subjects and smooth motions.
@@ 1 Introduction
Text-to-video (T2V) generation [1, 2, 3, 4, 5] has achieved fantastic progress taking advantages of diffusion models [6, 7, 8]. Recently, artists have dreamed of generating videos with their own belongings, e.g., pets, which directs a new research direction named customized T2V generation. Although existing methods [9, 10, 11] have been proposed to generate videos from a single object, tackling multiple objects still remains a difficult scenario. The key challenge is to ensure the co-occurrence of multiple objects in the generated video and retain their corresponding identities.
@@
A recent work, VideoDreamer [12], proposes disen-mix finetuning and human-in-the-loop re-finetuning strategies based on Stable Diffusion [13], aiming to generate videos from multiple subjects. However, VideoDreamer falls short in guaranteeing the co-occurrence of multiple subjects and disentangling similar subjects due to its inconsistent object mixing strategy. As shown in Figure 3, both the car and cat can not be consistently generated across all frames. Another straight forward solution for multi-subject driven T2V generation is by combining multi-subject driven Text-to-Image (T2I) and Image-to-Video (I2V) models. However, this naive approach fails from two aspects, including inaccurate T2I customization and lack of motion (see Figure 3 ‘DisenDiff [14] + SVD [15]’).
@@
In contrast, in our approach, we ensure the co-occurrence of multiple objects during model training, which encourages the model to capture the presence of different subjects simultaneously, thereby facilitating co-occurrence during inference. Additionally, we propose an attention control mechanism to disentangle multiple subjects during training, effectively guiding the model to focus on the corresponding subject area while disregarding irrelevant parts of the image. To facilitate this process, we incorporate a ground truth object mask, obtained through segmentation either from a model like SAM [16] or provided by human annotators, as supervision during optimization. Our attention mechanism consists of two key designs that contribute to the disentanglement of subjects. Firstly, we highlight the corresponding subject area using the ground truth object mask on the cross-attention map, promoting better preservation of subject identity. Secondly, we optimize the cross-attention map towards a slight negative value, excluding the desired subject and mitigating the influence of irrelevant area in the input image. To comprehensively evaluate our proposed approach, we have curated a diverse dataset covering a wide range of 13 categories. Beyond the MultiStudioBench [12] which includes 12 subject pairs, our dataset comprises 63 individual subjects and 68 subject pairs, notably featuring some challenging scenarios involving visually similar objects. Through extensive experiments on this benchmark dataset, we provide qualitative, quantitative, and user study results that demonstrate the superiority of our method in generating high-quality videos with customized subjects. In summary, our contributions are as follows:
@@
  • We collect a multi-subject T2V dataset and build it as a comprehensive benchmark, which covers a wide range of subject categories and diverse subject pairs over them.
  • Our method consistently outperforms previous state-of-the-art (SOTA) approaches in qualitative, quantitative and human preference evaluation, e.g., 11.99% CLIP Image Aligenment and 23.39% DINO Image Aligenment better than the previous best method.
@@ 2 Related Work
Text-to-Video Generation.
@@
Text-to-video generation has made significant advancements in recent years [17, 18, 1]. Early approaches [19, 20, 21, 22] employed GANs [23] and VQ-VAE [24], while more recent works have explored diffusion models to generate high-quality videos [25, 26, 27, 28]. Make-A-Video [29] utilizes a pre-trained image diffusion model with additional temporal attention finetuning. VideoLDM [30] introduces a multi-stage alignment approach in the latent space to generate high-resolution and temporally consistent videos. Other methods [31, 32] generate videos with an image as the first frame and randomly initialized subsequent frames. To enhance controllability, VideoComposer [33] incorporates additional guidance signals, such as depth maps, to produce desired videos alongside text inputs. VideoDirectorGPT [34] aims to control the video generation with the guidance from different scenes and specific layouts generated by GPT4 in the temporal axis. GEST [35] models the representation between the text and video via a graph of events in the spatial and temporal space, in which the video is generated following a timeline. Tune-A-Video [36] proposes a temporal self-attention module that fine-tunes a pre-trained image diffusion model, achieving successful generation of videos with specific text guidance. Furthermore, diffusion-based video-to-video editing approaches [37, 38] have also been proposed for practical usages.
Subject-driven Customization.
@@
There has been a growing interest in customizing pre-trained image and video diffusion models for personalized generation [39, 40, 41, 42, 43, 44, 11]. Customization involves generating images and videos with specific subjects, typically based on a few reference images. For image diffusion customization, Textual Inversion [39] represents a specific object as a learnable text token using only a few reference images. This learned text token can then be integrated into a sentence to generate personalized images during inference. Additionally, DreamBooth [45] fine-tunes the weights of the diffusion model to improve fidelity in image generation. Several works [46, 14, 47, 48] have explored personalized image diffusion with multiple subjects, focusing on parameter-efficient finetuning and text embedding learning. While there have been initial attempts to customize video diffusion, such as VideoAssembler [9], VideoBooth [10], and ID-Animator [49], which use reference images to personalize the video diffusion model while preserving subject identity, and DreamVideo [11], which decouples the learning process for subject and motion customization, these methods are designed for single objects and cannot handle multiple subjects when given. A recent work, VideoDreamer [12], proposes multi-subject driven video customization through disen-mix finetuning strategy with LoRA [50]. However, the generated videos do not guarantee the co-occurrence of multiple subjects, or disentanglement of different subjects. In this work, we propose a simple yet effective co-occurrence and attention mechanism that disentangles multiple subjects using masks as guidance while preserving the co-occurrence of subjects in the generated videos.
@@ 3.1 Preliminary: Text-to-Video
Video diffusion models (VDMs) [1, 18, 17] generate videos by gradually denoising a randomly sampled Gaussian noise $\epsilon$, following an iterative denoising process that resembles a reverse procedure of a fixed-length Markov Chain. This iterative denoising allows VDMs to capture the temporal dependencies presented in video data. To be specific, a video diffusion model $\theta$ predicts the added noise at each timestep $t$ given a text condition $c$, where $t\in\{1,2,\cdots,T\}$. The training objective for this process can be expressed as a reconstruction loss:
@@
where $\boldsymbol{z}\in\mathbb{R}^{B\times L\times H\times W\times D}$ is the latent code of the input video with batch size $B$, video length $L$, height $H$, width $W$, and latent dimension $D$. $\epsilon_{\theta}$ is the noise prediction from the model. $\mathcal{T}$ is a pre-trained text encoder. $\boldsymbol{z}_{t}$ is obtained by adding noise to the ground truth $\boldsymbol{z}_{0}$ with $\boldsymbol{z}_{t}=\alpha_{t}\boldsymbol{z}_{0}+\sqrt{1-\alpha_{t}^{2}}\epsilon$, where $\alpha_{t}$ is a diffusion hyperparameter. In this work, we utilize zeroscope [51] T2V model as our base model, which is built upon a 3D U-Net, with spatial and temporal modeling for generating high-quality videos.
@@ 3.2 CustomVideo with Multiple Subjects
An overview of our proposed CustomVideo framework is illustrated in Figure 2. For customization perspective, we introduce a new learnable word token for every subject, e.g., ‘<new1>’ for cat and ‘<new2>’ for dog, representing the corresponding identity. During the training stage, we employ a concatenation technique to ensure the co-occurrence of multiple subjects. Specifically, we combine these subjects into a single image, facilitating model’s learning of multi-subject patterns. To address the challenge of entanglement among highly similar subjects, we propose an attention control mechanism. This mechanism ensures that the learnable word tokens align with their corresponding regions on the cross-attention map (see Figure 6). By achieving this alignment, we enable the model to disentangle different subjects and enhance the quality of the generated videos. The training process focuses on training the subject-related learnable word tokens, as well as the key and value weights within the cross-attention layers of the U-Net architecture, adopting a parameter-efficient finetuning approach. During the inference stage, users only need to provide a text prompt description integrated with the corresponding learned word tokens to generate high-quality videos aligned with their preferences. In the following, we will delve into the detailed workings of CustomVideo.
@@
In the context of multi-subject driven T2V generation, ensuring that the model consistently learns to generate videos with multiple subjects is crucial. In a previous work, VideoDreamer [12], the authors proposed a disen-mix strategy for multi-subject driven generation by fine-tuning the model using both single subject images and concatenated images of multiple subjects. However, we have identified that this mixing strategy can confuse the model due to the inconsistent number of subjects present in a single image. Consequently, it leads to unstable generation of videos with multiple subjects (see line ‘both single and concat’ in Figure 5). In our approach, we have found that fine-tuning the model using only concatenated images of multiple subjects is sufficient to ensure the co-occurrence of multiple subjects in the generated videos. Additionally, we have observed that providing clear subjects without background aids the model in learning the specific characteristics of the subjects more effectively. To achieve this, we perform background removal on the subject images. This can be accomplished manually or by employing an automatic tool such as the SAM model [16].
Figure 2: The overview of our proposed CustomVideo. We propose a simple yet effective co-occurrence and attention control mechanism with mask guidance to preserve the fidelity of subjects for multi-subject driven text-to-video generation. During the training stage, only the key and value weights in the cross-attention layers are fine-tuned, together with a learnable word token for every subject. In the inference stage, given a text prompt integrated with learned word tokens, we can easily obtain high-quality videos with specific subjects.
@@
Ensuring the co-occurrence of multiple subjects in generated videos is achieved through the above. However, a more challenging task is preserving the distinct characteristics of each subject without interference. Simply fine-tuning the model with concatenated images can lead to confusion among subject characteristics. As shown in Figure 5 (line ‘w/o pos. attn.’), the generated cat predominantly resembles the shape of the provided dog sample, despite having similar color and texture to the provided cat. Thus, disentangling the multiple subjects becomes crucial to generate high-quality videos that faithfully represent each subject’s characteristics.
@@ 3.3 Model Training and Inference
During training, we only fine-tune the weights of key and value in all of the cross-attention layers. We use a more flexible approach by fine-tuning the model with subject images extended to single-frame videos as both input and supervision, rather than using subject videos. This approach guarantees the co-occurrence of multiple subjects because subject images are much easier to obtain and smoother to concatenate compared to subject videos. Moreover, fine-tuning with subject images saves lots of computation cost. Inspired by previous T2I personalization [45, 40], we conduct class-specific prior preservation to improve the diversity of generated videos and alleviate the issue of language drift. The loss for prior preservation is
@@
where $\boldsymbol{z}^{\prime}$ is the latent code of the input class image, $\epsilon_{\theta}^{\prime}$ is the noise prediction from the model $\theta$, $\epsilon^{\prime}$ is the randomly sampled Gaussian noise, $c^{\prime}$ is the text condition for class image, and $t^{\prime}$ is the sampling timestep. To this end, we train our CustomVideo via an end-to-end manner, with the following overall training objective:
@@
During inference, CustomVideo only requires a specific text prompt with corresponding learned word token integrated to generate a required video. To be note that ground truth masks of subjects are not required during inference. Although the model is fine-tuned with single-frame videos extended from subject images, our proposed CustomVideo can generate videos with high diversity while not losing the the fidelity of subjects and motion smoothness (see Figure 1 and Figure 3).